# Explicit Runge-Kutta Methods with MQ and IMQ-Radial Basis Functions

### v1

@@ 4 Convergence of RBF RUNGE-KUTTA Method
$k_{2}(w)=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{w}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right),$
$k_{3}(w)=f\left(t_{n}+c_{3}h,\sqrt{(1+\epsilon_{n3}^{2}a_{31}^{2}h^{2})}hk_{1}+\frac{w}{\sqrt{1+\epsilon_{n3}^{2}c_{2}^{2}h^{2}}}\right).$
@@
$\quad-f\left(t_{n}+c_{3}h,\ \sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\,\,h(a_{31}k_{1}(v_{n})+a_{32}k_{2}(v_{n}))+\frac{v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right)\Bigg{|}$
$\leq L\Bigg{|}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}h\left(a_{31}(k_{1}(u_{n})-k_{1}(v_{n}))+a_{32}(k_{2}(u_{n})-k_{2}(v_{n}))\right)+\frac{u_{n}-v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\Bigg{|}$
@@
$\quad\left.+\frac{L|u_{n}-v_{n}|}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}+\frac{|u_{n}-v_{n}|}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right].$
@@
$+b_{3}hL\Biggr{[}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{31}L+\sqrt{1+\epsilon_{n3}h^{2}c_{3}^{2}}\,\,ha_{32}\Biggr{(}L^{2}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{21}+\frac{L}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\Biggl{)}\Biggr{]}|u_{n}-v_{n}|$
@@
$\quad+b_{3}hL\left[\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{31}L+\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{32}\left(L^{2}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{21}+\frac{L}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right)\right]$
@@
$\left.+a_{32}b_{3}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\right)+b_{3}L^{3}h^{3}a_{21}a_{32}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}}$
@@
$\left.+a_{32}b_{3}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\right)+b_{3}L^{3}h^{3}a_{21}a_{32}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}}$
$\leq(1+\mathcal{C}_{3}h)e^{hL}$
$\leq e^{h(1+\mathcal{C}_{3})}$
@@
$\mathcal{C}_{3}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{\begin{aligned} &h^{2}L^{3}b_{3}c_{2}a_{32}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}-\frac{h^{2}L^{2}}{2}\biggl{(}b_{2}c_{2}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\\ &+b_{3}a_{31}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}+b_{3}a_{32}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\biggr{)}\end{aligned}}{e^{hL}},$
@@ 5 Stability regions
Stability regions indicates to the set of values in complex plane where absolute value of the stability polynomial is bounded by one. $y^{\prime}=\lambda y$ is taken as the test function. Substituting the test function in the expression of the method and replacing $\lambda h$ by $z$, $h$ is step length, we obtain ratio of $v_{n+1}$ and $v_{n}$. The ratio is termed as stability polynomial and it is in the form $v_{n+1}=R(z)v_{n}$, where $R(z)$ is stability polynomial. Stability region is defined by $|R(z)|\leq 1$. Bounding the stability polynomial by one ensures that the error will not grow unboundedly. Stability polynomials of the two, three, and four stage claasical Runge-Kutta methods are as follows
@@
$R(z)$ $=1+z+\frac{1}{2}z^{2}+\frac{1}{6}z^{3}+\frac{1}{24}z^{4},$ (5.3)
which in general indicate the absolute stability region. The stability polynomial for two stage MQ RBF Runge-Kutta method (3.1) is given by
@@
(a) Stability regions of RK2 (red) and IMQ-RK2 (blue).
(b) Stability regions of RK2 (red) and MQ-RK2 (blue).
(c) Stability regions of RK3 and IMQ-RK3.
(d) Stability regions of RK3 and MQ-RK3.
Figure 1: Stability region for various classical, MQ and IMQ-Runge Kutta schemes
The stability polynomial of the two stage IMQ RBF Runge-Kutta method (3.2) is
$R(z)=1+\frac{z}{4}+\frac{1}{2}z^{2}\sqrt{(1-\frac{4z^{2}}{9})}+\frac{3}{4}\frac{z}{\sqrt{(1-\frac{4z^{2}}{9}})}.$
The stability polynomial of the two-stage IMQ RBF Runge-Kutta method (3.5) are
@@ 6 Numerical results
In this section, numerical examples of MQ and IMQ RBF Runge-Kutta method are presented and discussed in support of the theory presented in previous sections. All calculations are performed with MATLAB R2022A. A comparison is followed with the corresponding classical Runge-Kutta method. All the errors preseneted here are calculated using $L_{1}-$norm.
@@
Table 2: Global error and order of accuracy for example 6.1 by RK2 and IMQ-RK2.
@@
$\frac{du}{dt}=-4t^{3}u^{2},\quad-10<t\leq 0,\quad u(-10)=\frac{1}{10001}$
This is a stiff problem with the exact solution $u(t)=\frac{1}{t^{4}+1}$ where the solution exhibits rapid variation over the interval $[-10,0]$. The global error is evaluated at $t=0$. Table 4 shows that the numerical solution achieves second-order accuracy with the standard RK2 method and third-order accuracy with the enhanced variants MQ-RK2 and IMQ-RK2. As shown in Table 5, the various RK3 and IMQ-RK3 methods yield third and fourth order accuracy, respectively. These RK3 and IMQ-RK3 methods are discussed in Sections 2 and 3, respectively. Table 6 presents the global error and confirms fourth-order accuracy for the ODE using various MQ-RBF methods described in Section 3.
@@
Table 5: Global error and order of accuracy for example 6.2 by RK2, IMQ-RK2.
@@
Consider,
$\frac{du}{dt}=\frac{2t^{2}-u}{t^{2}u-t},\quad 1<t\leq 2,\quad u(1)=2,$
@@
Table 8: Global error and order of accuracy for example 6.3 by RK and IMQ-RK3.
@@ 7 Conclusions
In this study, we introduced MQ-RBF and IMQ-RBF-based Runge-Kutta methods, which demonstrate improved accuracy orders compared to classical Runge-Kutta schemes. A detailed analysis of convergence behavior, stability regions, and local convergence order was conducted. The results indicate that although the IMQ-RBF method incurs higher computational complexity, it consistently outperforms the MQ-RBF method in terms of both numerical error and convergence order. However, implementing fourth-stage, fifth-order Runge-Kutta methods with MQ and IMQ radial basis functions (RBFs) on standard CPU architectures presents significant memory challenges, complicating the search for optimal shape parameters. To address these limitations, the development of memory-efficient MQ and IMQ RBF-based fourth-order, five-stage Runge-Kutta methods is currently underway and will be presented in future work.

### v2

@@ 4 Convergence of RBF Runge-Kutta Methods
$k_{2}(w)$ $=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}a_{21}^{2}h^{2})}\,ha_{21}k_{1}+\frac{w}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right),$
$k_{3}(w)$ $=f\left(t_{n}+c_{3}h,\sqrt{(1+\epsilon_{n3}^{2}c_{3}^{2}h^{2})}\,h(a_{31}k_{1}+a_{32}k_{2})+\frac{w}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right).$
@@
$\qquad-f\left(t_{n}+c_{3}h,\ \sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\,\,h(a_{31}k_{1}(v_{n})+a_{32}k_{2}(v_{n}))+\frac{v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right)\Bigg|,$
$\leq L\Bigg|\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}h\left(a_{31}(k_{1}(u_{n})-k_{1}(v_{n}))+a_{32}(k_{2}(u_{n})-k_{2}(v_{n}))\right)+\frac{u_{n}-v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\Bigg|,$
@@
$\quad\left.+\frac{L|u_{n}-v_{n}|}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}+\frac{|u_{n}-v_{n}|}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right],$
@@
$+b_{3}hL\Biggr[\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{31}L+\sqrt{1+\epsilon_{n3}h^{2}c_{3}^{2}}\,\,ha_{32}\Biggr(L^{2}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{21}+\frac{L}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\Biggl)\Biggr]|u_{n}-v_{n}|.$
@@
$\quad+b_{3}hL\left[\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{31}L+\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\,\,ha_{32}\left(L^{2}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\,\,ha_{21}+\frac{L}{\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}}\right)\right],$
@@
$\left.+a_{32}b_{3}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}a_{21}^{2}h^{2}}\right)+b_{3}L^{3}h^{3}a_{21}a_{32}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}},$
@@
$\left.+a_{32}b_{3}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\right)+b_{3}L^{3}h^{3}a_{21}a_{32}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}\sqrt{1+\epsilon_{n2}^{2}h^{2}a_{21}^{2}},$
$\leq(1+\mathcal{C}_{3}h)e^{hL},$
$\leq e^{h(1+\mathcal{C}_{3})},$
@@
$\mathcal{C}_{3}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{A}{e^{hL}},$
with
$A=$ $h^{2}L^{3}b_{3}c_{2}a_{32}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}$
$\qquad-\frac{h^{2}L^{2}}{2}\biggl(b_{2}c_{2}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}a_{31}\sqrt{1+\epsilon_{n3}^{2}h^{2}c_{3}^{2}}+b_{3}a_{32}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\biggr).$
@@ 5 Stability Regions
Stability regions indicates to the set of values in complex plane where absolute value of the stability polynomial is bounded by one. $y^{\prime}=\lambda y$ is taken as the test function. Substituting the test function in the expression of the method and replacing $\lambda h$ by $z$, $h$ is step length, we obtain ratio of $v_{n+1}$ and $v_{n}$. The ratio is termed as stability polynomial and it is in the form $v_{n+1}=R(z)v_{n}$, where $R(z)$ is stability polynomial. Stability region is defined by $|R(z)|\leq 1$. Bounding the stability polynomial by one ensures that the error will not grow unboundedly. Stability polynomials of the two, three, and four stage classical Runge-Kutta methods are as follows
@@
which in general indicate the absolute stability region. The stability polynomial for two stage MQ-RBF Runge-Kutta method (3.3) is given by
@@
The stability polynomial of the two stage IMQ-RBF Runge-Kutta method (3.4) is
$R(z)=1+\frac{z}{4}+\frac{1}{2}z^{2}\sqrt{(1-\frac{4z^{2}}{9})}+\frac{3}{4}\dfrac{z}{\sqrt{(1-\frac{4z^{2}}{9}})}.$
The stability polynomial of the three-stage IMQ-RBF Runge-Kutta method (3.8) are
@@
(a) Stability regions of RK2 and IMQ-RK2.
(b) Stability regions of RK2 and MQ-RK2.
(c) Stability regions of RK3 and IMQ-RK3.
(d) Stability regions of RK3 and MQ-RK3.
Figure 1: Stability region for various classical, MQ and IMQ -RBF Runge Kutta schemes
@@ 6 Numerical Results
In this section, numerical examples of MQ and IMQ RBF Runge-Kutta methods are presented and discussed in support of the theory presented in previous sections. All calculations are performed with MATLAB R2022A. A comparison is followed with the corresponding classical Runge-Kutta method. All the errors preseneted here are calculated using $L_{1}$-norm.
@@
Table 2: Global error and order of accuracy for example 6.1 by RK3 and IMQ-RK3.
@@
$\frac{du}{dt}=-4t^{3}u^{2},\quad-10<t\leq 0,\quad u(-10)=\frac{1}{10001}.$
This is a stiff problem with the exact solution $u(t)=\frac{1}{t^{4}+1}$ where the solution exhibits rapid variation over the interval [10,0][-10,0]. The global error is evaluated at $t=0$. Table 4 shows that the numerical solution achieves second-order accuracy with the standard RK2 method and third-order accuracy with the enhanced variants MQ-RK2 and IMQ-RK2. As shown in Table 5, the various RK3 and IMQ-RK3 methods yield third and fourth order accuracy, respectively. These RK3 and IMQ-RK3 methods are discussed in Sections 2 and 3, respectively. Table 6 presents the global error and confirms fourth-order accuracy for the ODE using various MQ-RBF methods described in Section 3.
@@
Table 5: Global error and order of accuracy for example 6.2 by RK3, IMQ-RK3.
@@
Consider
$\frac{du}{dt}=\frac{2t^{2}-u}{t^{2}u-t},\quad 1<t\leq 2,\quad u(1)=2.$
@@
Table 8: Global error and order of accuracy for example 6.3 by RK3 and IMQ-RK3.
@@
Example 6.4
Explicit RBF-Runge method for scalar is extended to system case by using component wise operation. We consider the system of equations
$\frac{du_{1}}{dt}$ $=e^{t}-5u_{1}+3u_{2},$
$\frac{du_{2}}{dt}$ $=-3u_{1}+u_{2},$
with initial data
$u_{1}(0)=1,u_{2}(0)=0.$
The exact solution of the above ordinary differential equation $u_{1}(t)=(1-2t)e^{2}t,\,\,u_{2}(t)=(\frac{1}{3}-2t)e^{-2t}-\frac{1}{3}e^{t}$. Solution is approximated at final time $t=5$ and shown in table 10. It is observed that truncation error appeared in RK-2 method follows second order accuracy, whereas MQ-RK2 and IMQ-RK2 attains third order accuracy. Furthermore, the error decreases as number of grid points increases.
Table 10: Global error and order of accuracy for example 6.4 by RK-2, MQ-RK2, IMQ-RK2.
$u_{1}$ RK 2 IMQ-RK2 MQ-RK2
N Error Order Error Order Error Order
20 3.8578e-01 3.0277e-02 4.0712e-02
40 6.6058e-02 2.5460 6.3264e-03 2.2588 3.2458e-03 3.6488
80 1.3425e-02 2.2989 7.1610e-04 3.1431 3.4340e-04 3.2406
160 3.0077e-03 2.1582 8.2946e-05 3.0887 4.0364e-05 3.0887
320 7.1054e-04 2.0817 9.5653e-06 3.1163 5.2049e-06 2.9551
640 1.7259e-04 2.0415 1.0391e-06 3.2024 7.7303e-07 2.7513
$u_{2}$ RK 2 IMQ-RK2 MQ-RK2
N Error Order Error Order Error Order
20 2.6347e-02 2.8261e-02 3.6335e-03
40 2.7899e-02 -0.0826 2.9749e-03 3.2479 1.8151e-03 1.0013
80 9.0268e-03 1.6279 4.6315e-04 2.6833 3.6812e-05 2.7065
160 2.4810e-03 1.8633 6.4613e-05 2.8416 2.7808e-04 2.9173
320 6.4646e-04 1.9403 8.8737e-06 2.8642 4.3812e-06 3.0708
640 1.6478e-04 1.9720 1.2748e-06 2.7992 4.1112e-07 3.4137
Example 6.5
The Duffing equation [7] is given by
$q^{\prime\prime}+\omega^{2}q=k^{2}\left(2q^{3}-q\right),0\leq t\leq 20,$
$q(0)=0,q^{\prime}(0)=\omega,$
with $0\leq k<\omega$. Choosing $p=q^{\prime}$ and $u=(p,q)^{T}$, the above-mentioned second-order differential equation reduces to a first-order differential equation system in the form $u^{\prime}=f(u)$, where $f(u)=(-\omega^{2}q+k^{2}(2q^{3}-q),\,p)^{T}$. The analytic solution is given by $q(t)=sn(\omega,k/\omega).$ Here, $\omega=10$, $k=0.03$ and $h=0.05$ is taken in our experimental results. From figure 2, it is clearly visible that the exact solution blow up over time whereas, the approximate solution computed by MQ-RK2 and IMQ-RK2 method damp over time. Duffing equation can be written as Hamiltonian system, so it is structure preserving but RK2, MQ-RK2, IMQ-RK2 are not so. So, while approximating the Duffing equation it is either blowing up or damping. In our future work, we will explore the sympletic Runge-Kutta method.
Figure 2: The three graphs represents solution of example 5 where approximate solution by RK2, MQ-RK2 and IMQ-RK2 are compared with exact solution respectively at final time $t=20$.
@@ 7 Conclusions
In this study, we have introduced MQ-RBF and IMQ-RBF-based Runge-Kutta methods, which demonstrate improved accuracy orders compared to classical Runge-Kutta schemes. A detailed analysis of the convergence behavior, stability regions, and local convergence order was performed through five numerical tests. The results indicate that although the IMQ-RBF method incurs higher computational complexity, it consistently outperforms the MQ-RBF method in terms of both numerical error and convergence order. Furthermore, implementing fourth-stage, fifth-order Runge-Kutta methods with MQ and IMQ radial basis functions (RBFs) on standard CPU architectures presents significant memory challenges, complicating the search for optimal shape parameters. To address these limitations, the development of memory-efficient MQ and IMQ RBF-based fourth-order, five-stage Runge-Kutta methods is currently underway and will be presented in future work.
Acknowledgments The author Shipra Mahata is supported by UGC, India (NTA Ref No. 211610046691).